An expandable container of oxygen gas has a volume of 12.5mL at a temperature of 26.0°C. What volume will the gas occupy at 54.0°C?


1
Expert's answer
2022-05-18T15:02:03-0400

One of the forms of Charles’s low when the pressure is constant:

"\frac{V_1}{T_1} = \frac{V_2}{T_2}"

With the increase of temperature volume of a gas become:

"V_2 = \frac{V_1 \times T_2}{T_1} = \frac{12.5 ml \times 54}{26} = 26 ml"